# Approximating Two-Layer ReLU Networks for Hidden State Analysis in Differential Privacy

### v2

@@ 3 Convex Approximation of Two-Layer ReLU Networks
We next derive step by step the strongly convex approximation of the 2-layer ReLU minimization problem and show that the derived problem is amenable for the privacy amplification by iteration type of $(\varepsilon,\delta)$-DP analysis. To simplify the presentation, we consider a 1-dimensional output network (e.g., a binary classifier). It will be straightforward to construct multivariate output networks from the scalar networks (see also Ergen et al.,, 2023).
@@ 3.1 Convex Duality of Two-layer ReLU Problem
We first consider the convex reformulation of the 2-layer ReLU minimization problem as presented by Pilanci and Ergen (2020). In particular, consider training a 2-layer ReLU network (with hidden-width $m$) $f\,:\,\mathbb{R}^{d}\rightarrow\mathbb{R}$,
$f(x)=\sum\nolimits_{j=1}^{m}\phi(u_{j}^{T}x)\alpha_{j},$ (3.1)
where the weights $u_{i}\in\mathbb{R}^{d}$, $i\in[m]$ and $\alpha\in\mathbb{R}^{m}$, and $\phi$ is the ReLU activation function, i.e., $\phi(t)=\max\{0,t\}$. For a vector $x$, $\phi$ is applied element-wise, i.e. $\phi(x)_{i}=\phi(x_{i})$.
Using the squared loss and $L_{2}$-regularization with a regularization constant $\lambda>0$, the 2-layer ReLU minimization problem can then be written as
@@
The convex reformulation of this problem is based on enumerating all the possible activation patterns of $\phi(Xu)$, $u\in\mathbb{R}^{d}$. The set of activation patterns that a ReLU output $\phi(Xu)$ can take for a data feature matrix $X\in\mathbb{R}^{n\times d}$ is described by the set of diagonal boolean matrices
$\mathcal{D}_{X}=\{D=\mathrm{diag}(\mathds{1}(Xu\geq 0))\,:\,u\in\mathbb{R}^{d}\},$
where for $i\in[n]$, $\big{(}\mathds{1}(Xu\geq 0)\big{)}_{i}=1$, if $(Xu)_{i}\geq 0$ and 0 otherwise. Here $\left|\mathcal{D}_{X}\right|$ is the number of regions in a partition of $\mathbb{R}^{d}$ by hyperplanes that pass through origin and are perpendicular to the rows of $X$. We have (Pilanci and Ergen,, 2020):
$\left|\mathcal{D}_{X}\right|\leq 2r\left(\frac{{\rm e}\hskip 1.0pt(n-1)}{r}\right)^{r},$
where $r=\mathrm{rank}(X)$.
Let $\left|\mathcal{D}_{X}\right|=M$ and denote $\mathcal{D}_{X}=\{D_{1},\ldots,D_{M}\}$. Let $\lambda>0$. Next, we the parameter space is partitioned into convex cones $C_{1},\ldots,C_{M}$, $C_{i}=\{u\in\mathbb{R}^{d}\,:\,(2D_{i}-I)Xu\geq 0\}$, and we consider a convex optimization problem with group $\ell_{2}$ $\ell_{1}$ regularization
$\min_{v_{i},w_{i}}\frac{1}{2}\left\lVert\sum\nolimits_{i\in[M]}D_{i}X(v_{i}-w_{i})-y\right\rVert_{2}^{2}+\lambda\sum\nolimits_{i\in[M]}(\left\lVert v_{i}\right\rVert_{2}+\left\lVert w_{i}\right\rVert_{2})$ (3.3)
such that for all $i$, $1\leq i\leq M:\,v_{i},w_{i}\in C_{i}$ i.e.
$(2D_{i}-I)Xw_{i}\geq 0,\quad(2D_{i}-I)Xv_{i}\geq 0.$
Interestingly, for a sufficiently large hidden-width $m$, the ReLU minimization problem (3.2) and the convex problem (3.3) have equal minima.
Theorem 3.1 (Pilanci and Ergen, 2020, Thm. 1).
There exists $m^{*}\in\mathbb{N}$, $m^{*}\leq d+1$, such that for all $m\geq m^{*}$, the ReLU minimization problem (3.2) and the convex problem (3.3) have equal minima.
Moreover, Pilanci and Ergen, (2020) show that for a large enough hidden-width $m$, the optimal weights of the ReLU network can be constructed from the optimal solution of the convex problem 3.3. Subsequent work, such as (Mishkin et al.,, 2022), extends the equivalence in Thm. 3.1 to general convex loss functions $\mathcal{L}$, rather than focusing solely on the squared loss. For simplicity, we focus on the squared loss in our presentation. We remark that convex formulations have also be shown for two-layer convolutional networks (Bartan and Pilanci,, 2019) and for multi-layer ReLU networks (Ergen and Pilanci,, 2021).
@@ 3.2 Stochastic Approximation
Since $\left|\mathcal{D}_{X}\right|$ is generally an enormous number, stochastic approximations to the problem (3.3) have been considered  (Pilanci and Ergen,, 2020; Wang et al.,, 2022; Mishkin et al.,, 2022; Kim and Pilanci,, 2024). In this approximation, vectors $u_{i}\sim\mathcal{N}(0,I_{d})$, $i\in[P]$, $P\ll M$, are sampled randomly to construct the boolean diagonal matrices $D_{1},\ldots,D_{P}$, $D_{i}=\mathrm{diag}(\mathds{1}(Xu_{i}\geq 0))$, and the problem (3.3) is replaced by
$\min_{v_{i},w_{i}}\,\,\,\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}D_{i}X(v_{i}-w_{i}),y\right\rVert_{2}^{2}+\lambda\sum\nolimits_{i=1}^{P}(\left\lVert v_{i}\right\rVert_{2}+\left\lVert w_{i}\right\rVert_{2})$ (3.4)
such that for all $i\in[P]$ : $v_{i},w_{i}\in C_{i}$, i.e.,
$(2D_{i}-I)Xw_{i}\geq 0,\quad\quad(2D_{i}-I)Xv_{i}\geq 0.$ (3.5)
For practical purposes we consider a stochastic approximation of this kind. However, the constraints 3.5 are data-dependent which potentially makes private learning of the problem (3.4) difficult. Moreover, the overall loss function given by Eq. 3.4 is not generally strongly convex which prevents us using privacy amplification results such as Theorem 2.8 for NoisyCGD. We next consider a strongly convex problem without constraints of the form 3.5.
@@ 3.3 Stochastic Strongly Convex Approximation
Motivated by experimental observations and also the formulation given in (Wang et al.,, 2022), we consider global minimization of the loss function (denote $v=\{v_{i}\}_{i=1}^{P}$)
$\mathcal{L}\big{(}v,X,y\big{)}=\frac{1}{2n}\left\lVert\sum\nolimits_{i=1}^{P}D_{i}Xv_{i}-y\right\rVert_{2}^{2}+\frac{\lambda}{2}\sum\nolimits_{i=1}^{P}\left\lVert v_{i}\right\rVert_{2}^{2},$ (3.6)
where the diagonal boolean matrices $D_{1},\ldots,D_{P}\in\mathbb{R}^{n\times n}$ are constructed by taking first $P$ i.i.d. samples $u_{1},\ldots,u_{P}$, $u_{i}\sim\mathcal{N}(0,I_{d})$, and then setting the diagonal elements of $D_{i}$’s as above as
$(D_{i})_{jj}=\max\big{(}0,\mathrm{sign}(x_{j}^{T}u_{i})\big{)}.$
Note that we may also write the loss function of Eq. (3.6) in the summative form
@@
where
$\ell(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(D_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}+\frac{\lambda}{2}\sum\nolimits_{i=1}^{P}\left\lVert v_{i}\right\rVert_{2}^{2},\quad(D_{i})_{jj}=\mathds{1}(x_{j}^{T}u_{i}\geq 0).$ (3.8)
Inference Time Model.
At the inference time, having a data sample $x\in\mathbb{R}^{d}$, using the $P$ vectors $u_{1},\ldots,u_{P}$ that were used for constructing the boolean diagonal matrices $D_{i}$, $i\in[P]$, used in the training, the prediction is carried out similarly using the function
$g(x,v)=\sum\nolimits_{i=1}^{P}\mathds{1}(u_{i}^{T}x\geq 0)\cdot x^{T}v_{i}$
@@
In experiments, we use cross-entropy loss instead of the mean square loss for the loss functions $\mathcal{L}_{j}$. Above, we have considered scalar output networks. In case of $k$-dimensional outputs and $k$-dimensional labels, we will simply use $k$ independent linear models parallely meaning that the overall model has a dimension $d\times P\times k$, where $d$ is the feature dimension and $P$ the number of randomly chosen hyperplanes.
@@ 3.4 Meeting the Requirements of DP Analysis
From Eq. 3.7 and Eq. (3.8) it is evident that each loss function $\ell(v,x_{j},y_{j})$, $j\in[n]$, depends only on the data entry $(x_{j},y_{j})$. In case $\lambda=0$, we can apply DP-SGD by simply clipping the sample-wise gradients $\nabla_{v}\ell(v,x_{j},y_{j})$. Moreover, in case $\lambda>0$, by clipping the data sample-wise gradients $\nabla_{v}h(v,x_{j},y_{j})$, where $h(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(D_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}$, the loss function $\ell$ becomes $2L$-sensitive (see Def. 2.7).
We also need to analyse the convexity properties of the problem (3.6) for the DP analysis of NoisyCGD. We have the following Lipschitz-bound for the gradients of the loss function (3.6).
@@
Due to the $L_{2}$-regularization, the loss function of the problem (3.6) is clearly $\lambda$-strongly convex. The properties of $\lambda$-strong convexity and $\beta$-smoothness are preserved when clipping the gradients, see, e.g., Section E.2 of (Redberg et al.,, 2024) or Section F.4 of (Ye and Shokri,, 2022). Thus, the DP accounting for NoisyCGD using Thm. 2.8 is applicable also when clipping the gradients.
4 Theoretical Analysis: Utility Bounds in the Random Data Model
Using classical results from private empirical risk minimization (ERM) we illustrate the improved convergence rate when compared to private training of 2-layer ReLU networks. We emphasize that a rigorous analysis would require having a priori bounds for the gradient norms. In the future work, it will be interesting to find our whether techniques used for the private linear regression (see e.g., Liu et al.,, 2023; Avella-Medina et al.,, 2023; Varshney et al.,, 2022; Cai et al.,, 2021) could be used to get rid of the assumption on bounded gradients.
We consider for the problem (3.6) utility bounds with random data. This data model is also commonly used in the analysis of private linear regression (see, e.g., Varshney et al.,, 2022).
Recently, Kim and Pilanci, (2024) have given several results for convex problem (3.6) under the assumption of random data, i.e., when $X_{ij}\sim\mathcal{N}(0,1)$ i.i.d. Their results essentially tell that taking $d$ and $n$ large enough (s.t. $n\geq d$), we have that with $P={O}(\tfrac{n\log n/\gamma}{d})$ random hyperplane arrangements we get zero global optimum for the stochastic problem (3.6) with probability at least $1-\gamma-\tfrac{1}{(2n)^{8}}$. If we choose $P={O}\left(\tfrac{n\log n/\gamma}{d}\right)$ hyperplane arrangements, we have an ambient dimension $p=d\cdot P={O}(n\log\tfrac{n}{\gamma})$ and directly get the following corollary of Thm. D.3.
@@ 4 Theoretical Analysis: Utility Bounds in the Random Data Model
Consider applying the private gradient descent (Alg. 1) to the practical stochastic strongly convex problem (3.6) and assume the gradients stay bounded by a constant $L>0$. For any $\gamma>0$, there exists $d_{1}$ such that for all $d\geq d_{1}$, $n\geq d$, with probability at least $1-\gamma-\frac{1}{(2n)^{8}}$,
$\mathcal{L}(\theta^{\textrm{priv}},D)\leq\widetilde{O}\left(\frac{1}{\sqrt{n}\varepsilon}\right),$
where $\widetilde{O}$ omits the logarithmic factors.
5 DP Hyperparameter Tuning for DP-SGD and NoisyCGD
When comparing experimentally DP optimization methods, it is crucial to take into account the effect of the hyperparameter tuning on the privacy costs. The most relevant to our work are the randomized tuning methods given by Papernot and Steinke, (2022) and the subsequent privacy profile-based analysis by Koskela et al., (2024). These results hold for a tuning algorithm that outputs the best model of the $K$ alternatives, where $K$ is a random variable. We consider the case $K$ is Poisson distributed, however mention that also other alternatives exist that allow adjusting the balance between compute cost of training, privacy and accuracy. The DP bounds in this case can be described as follows. Let $Q(y)$ the density function of the quality score of the base mechanism ($y$ denoting the score) and $A(y)$ the density function of the tuning algorithm that outputs the best model of the $K$ alternatives. Let $A$ and $A^{\prime}$ denote the output distributions of the tuning algorithm evaluated on neighboring datasets $D$ and $D^{\prime}$, respectively. Then, the hockey-stick divergence between $A$ and $A^{\prime}$ can be bounded using the following result.
Theorem 5.1.
$H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(A||A^{\prime})\leq m\cdot\delta(\widehat{\varepsilon}),$ (5.1)
Theorem 5.1 says that the hyperparameter tuning algorithm is $\big{(}\widehat{\varepsilon},m\cdot\delta(\widehat{\varepsilon})\big{)}$-DP in case the base mechanism is $(\varepsilon_{1},\delta(\varepsilon_{1})$-DP. If we can evaluate the privacy profile for different values of $\varepsilon$, we can also optimize the upper bound (5.1). When comparing DP-SGD and NoisyCGD, we use the fact that DP-SGD privacy profiles are approximately those of the Gaussian mechanism for large compositions (Sommer et al.,, 2019) as follows. Suppose DP-SGD is $(\varepsilon^{*},\delta^{*})$-DP for some values of the batch size $b$ noise scale $\sigma$ and number of iterations $T$. Then, we fix the value of the constant $c=\eta\cdot\lambda$ for NoisyCGD such that it is also $(\varepsilon^{*},\delta^{*})$-DP for the same hyperparameter values $b$, $\sigma$ and $T$, giving some GDP parameter $\mu^{*}$. Along the GDP privacy profile determined by $\mu^{*}$, we find $(\varepsilon_{1},\delta(\varepsilon_{1}))$ that optimizes the bound of Eq. (5.1) , and then evaluate the DP-SGD-$\delta$ using that same value $\varepsilon_{1}$, giving some value $\widehat{\delta}(\varepsilon_{1})$. Taking the maximum of $\delta(\varepsilon_{1})$ and $\widehat{\delta}(\varepsilon_{1})$ for the evaluation of $\widehat{\varepsilon}$ in the bound of Eq. (5.1) will then give a privacy profile that bounds the DP-guarantees of the hyperparameter tuning of both DP-SGD and NoisyCGD.

### v3

@@ 3 Convex Approximation of Two-Layer ReLU Networks
This section derives the strongly convex approximation of the 2-layer ReLU minimization problem and shows that the derived problem is amenable to privacy amplification by iteration analysis. Without loss of generality, we consider a 1-dimensional output network (e.g., a binary classifier), which can be extended to multivariate output networks later (see also Ergen et al., , 2023).
@@ 3.1 Convex Duality of Two-layer ReLU Problem
Consider training a ReLU network (with hidden-width $m$) $f\,:\,\mathbb{R}^{d}\rightarrow\mathbb{R}$ (Pilanci and Ergen, , 2020),
$f(x)=\sum\nolimits_{j=1}^{m}\phi(u_{j}^{T}x)\alpha_{j}.$ (3.1)
The weights are $u_{i}\in\mathbb{R}^{d}$, $i\in[m]$ and $\alpha\in\mathbb{R}^{m}$. The ReLU activation function is $\phi(t)=\max\{0,t\}$. For a vector $x$, $\phi$ is applied element-wise, i.e. $\phi(x)_{i}=\phi(x_{i})$.
Suppose the dataset $D$ consists of $n$ tuples of the form $z_{i}=(x_{i},y_{i})$, $x_{i}\in\mathbb{R}^{d}$, $y_{i}\in\mathbb{R}$, for $i\in[n]$. Using the squared loss and $L_{2}$-regularization with a regularization constant $\lambda>0$, the 2-layer ReLU minimization problem can be written as
@@
The convex reformulation is based on enumerating all the possible activation patterns of $\phi(Xu)$, $u\in\mathbb{R}^{d}$. The set of activation patterns that a ReLU output $\phi(Xu)$ can take for a data feature matrix $X\in\mathbb{R}^{n\times d}$ is described by the set of diagonal boolean matrices
$\mathcal{D}_{X}=\{\Lambda=\mathrm{diag}(\mathds{1}(Xu\geq 0))\,:\,u\in\mathbb{R}^{d}\},$ (3.3)
where for $i\in[n]$, $\big{(}\mathds{1}(Xu\geq 0)\big{)}_{i}=1$, if $(Xu)_{i}\geq 0$ and 0 otherwise. The number of regions in a partition of $\mathbb{R}^{d}$ by hyperplanes that pass through the origin and are perpendicular to the rows of $X$ is $\left|\mathcal{D}_{X}\right|$. We have by Pilanci and Ergen, (2020): $\left|\mathcal{D}_{X}\right|\leq 2r\left(\tfrac{{\rm e}\hskip 1.0pt(n-1)}{r}\right)^{r}$, where $r=\mathrm{rank}(X)$.
Let the number of activation patterns $\left|\mathcal{D}_{X}\right|=M$ and denote $\mathcal{D}_{X}=\{\Lambda_{1},\ldots,\Lambda_{M}\}$. Let $\lambda>0$. The parameter space is then partitioned into convex cones $\mathcal{V}_{1},\ldots,\mathcal{V}_{M}$, $\mathcal{V}_{i}=\{u\in\mathbb{R}^{d}\,:\,(2\Lambda_{i}-I)Xu\geq 0\}$, and we consider a convex optimization problem with group $\ell_{2}$ $\ell_{1}$ regularization
$\min_{v_{i},w_{i}}\frac{1}{2}\left\lVert\sum\nolimits_{i\in[M]}\Lambda_{i}X(v_{i}-w_{i})-y\right\rVert_{2}^{2}+\lambda\sum\nolimits_{i\in[M]}(\left\lVert v_{i}\right\rVert_{2}+\left\lVert w_{i}\right\rVert_{2})$ (3.4)
such that for all $i\in[M]:\,v_{i},w_{i}\in\mathcal{V}_{i}$ i.e., $(2\Lambda_{i}-I)Xw_{i}$ and $(2\Lambda_{i}-I)Xv_{i}\geq 0$ for all $i\in[M]$. Interestingly, for a sufficiently large hidden-width $m$, the ReLU minimization problem (3.2) and the convex problem (3.4) have equal minima (c.f., Thm.1 of Pilanci and Ergen, , 2020).
Moreover, Pilanci and Ergen, (2020) show that for a large enough hidden-width $m$, the optimal weights of the ReLU network can be constructed from the optimal solution of the convex problem 3.4. Subsequent work, such as (Mishkin et al., , 2022), extends the equivalence to general convex loss functions $\mathcal{L}$. For simplicity, we focus on the squared loss in our presentation. We remark that convex formulations have also been shown for two-layer convolutional networks (Bartan and Pilanci, , 2019) and multi-layer ReLU networks (Ergen and Pilanci, , 2021).
@@ 3.2 Stochastic Approximation
Since $\left|\mathcal{D}_{X}\right|$ is generally an enormous number, stochastic approximations to the problem (3.4) have been considered (Pilanci and Ergen, , 2020; Wang et al., , 2022; Mishkin et al., , 2022; Kim and Pilanci, , 2024). Vectors $u_{i}\sim\mathcal{N}(0,I_{d})$, $i\in[P]$, $P\ll M$, are then sampled randomly to construct the boolean diagonal matrices $\Lambda_{1},\ldots,\Lambda_{P}$, $\Lambda_{i}=\mathrm{diag}(\mathds{1}(Xu_{i}\geq 0))$, and the problem (3.4) is replaced by
$\min_{v_{i},w_{i}}\,\,\,\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}\Lambda_{i}X(v_{i}-w_{i}),y\right\rVert_{2}^{2}+\lambda\sum\nolimits_{i=1}^{P}(\left\lVert v_{i}\right\rVert_{2}+\left\lVert w_{i}\right\rVert_{2})$ (3.5)
such that for all $i\in[P]$ : $v_{i},w_{i}\in\mathcal{V}_{i}$, i.e., $(2\Lambda_{i}-I)Xw_{i}\geq 0$, and $(2\Lambda_{i}-I)Xv_{i}\geq 0$ for all $i\in[P]$. These constraints, however, are data-dependent, which makes private learning of the problem (3.5) difficult. Moreover, the overall loss function given by Eq. 3.5 is not generally strongly convex. This prevents using privacy amplification results such as Theorem 2.7 for NoisyCGD. We thus consider a strongly convex problem without any constraints and with a squared regularization term.
@@ 3.3 Stochastic Strongly Convex Approximation
Motivated by the needs of the DP analysis and the formulation given in (Wang et al., , 2022), we consider global minimization of the loss function
@@
with
$\ell(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(\Lambda_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}+\frac{\lambda}{2}\sum\nolimits_{i=1}^{P}\left\lVert v_{i}\right\rVert_{2}^{2},\quad(\Lambda_{i})_{jj}=\mathds{1}(x_{j}^{T}u_{i}\geq 0),$ (3.7)
where $v=\{v_{i}\}_{i=1}^{P}$, $v_{i}\in\mathbb{R}^{d}$ for $i\in[P]$ denote the learnable parameters. We note that due to the convention of the literature, the regularization is added term-wise (Bok et al., , 2024). The diagonal boolean matrices $\Lambda_{1},\ldots,\Lambda_{P}\in\mathbb{R}^{n\times n}$ are again constructed by taking first $P$ i.i.d. samples $u_{1},\ldots,u_{P}$, $u_{i}\sim\mathcal{N}(0,I_{d})$, and then setting the diagonals of $\Lambda_{i}$’s: $(\Lambda_{i})_{jj}=\max\big{(}0,\mathrm{sign}(x_{j}^{T}u_{i})\big{)}.$
The Model at Inference Time.
Having a sample $x\in\mathbb{R}^{d}$ at inference time, one uses the $P$ vectors $u_{1},\ldots,u_{P}$ that were used for constructing the boolean diagonal matrices $\Lambda_{i}$, $i\in[P]$ during training. The prediction is carried out similarly using the function $g(x,v)=\sum\nolimits_{i=1}^{P}\mathds{1}(u_{i}^{T}x\geq 0)\cdot x^{T}v_{i}.$
@@
In experiments, we use cross-entropy loss instead of the mean square loss for the loss functions $\mathcal{L}_{j}$. In case of $k$-dimensional outputs and $k$-dimensional labels, we use $k$ independent linear models in parallel. This means that the overall model has a dimension $d\times P\times k$, where $d$ is the feature dimension and $P$ the number of randomly chosen hyperplanes.
@@ 3.4 Meeting the Requirements of DP Analysis
In Eq. (3.7) each loss function $\ell(v,x_{j},y_{j})$, $j\in[n]$, depends only on the data entry $(x_{j},y_{j})$. By clipping the data sample-wise gradients $\nabla_{v}h(v,x_{j},y_{j})$, where $h(v,x_{j},y_{j})=\frac{1}{2}\left\lVert\sum\nolimits_{i=1}^{P}(\Lambda_{i})_{jj}x_{j}^{T}v_{i}-y_{j}\right\rVert_{2}^{2}$, the loss function $\ell$ becomes $2L$-sensitive (see Def. 2.6). Appendix B shows that the loss function $\ell(v,x_{j},y_{j})$ is a loss function of a generalized linear model and thus we are allowed to use the analysis of Bok et al., (2024) when clipping the gradients since then the clipped gradients are gradients of another convex loss (Song et al., , 2021). For the DP analysis, we must also analyze the loss function’s convexity properties (3.7).
We have the following Lipschitz bound for the gradients.
@@
Due to $L_{2}$-regularization, the loss function (3.7) is $\lambda$-strongly convex. The properties of $\lambda$-strong convexity and $\beta$-smoothness are preserved when clipping the sample-wise gradients $\nabla_{v}h(v,x_{j},y_{j})$ (Section E.2, Redberg et al., , 2024). Therefore, the DP accounting Thm. 2.7 is applicable with the same convexity parameters when clipping the gradients $\nabla_{v}h(v,x_{j},y_{j})$.
4 Theoretical Analysis
From a theoretical standpoint, convex models are advantageous over non-convex ones in private optimization. State-of-the-art empirical risk minimization (ERM) bounds for private convex optimization are of the order $O(\tfrac{1}{\sqrt{n}}+\tfrac{\sqrt{d}}{\varepsilon n})$, where $n$ is the number of training data entries, $d$ the dimension of the parameter space and $\varepsilon$ the DP parameter (Bassily et al., , 2019). In contrast, the bounds for non-convex optimization, which focus on finding stability points, are much worse, such as $O(\tfrac{1}{n^{1/3}}+\tfrac{d^{1/5}}{(\varepsilon n)^{2/5}})$ (Bassily et al., , 2021) and $O\left(\tfrac{d^{1/3}}{(\varepsilon n)^{2/3}}\right)$ (Arora et al., , 2023; Lowy et al., , 2024).
We consider utility bounds with random data for the problem (3.7). This data model is also commonly used in the analysis of private linear regression (see, e.g., Varshney et al., , 2022). Leveraging classical DP-ERM results, we derive utility bounds that are tighter than those obtainable for 2-layer ReLU networks via non-convex optimization. In addition to the convexity benefits of DP-ERM, we utilize the approximability of our convex models: with enough random hyperplanes, the global minimum $\mathcal{L}(\theta^{*},D)$ goes to zero with high probability.
4.1 Utility Bound for the Convex ReLU Approximation in the Random Data Model
Recently, Kim and Pilanci, (2024) have given several results for convex problems under the assumption of random data, i.e., when $X_{ij}\sim\mathcal{N}(0,1)$ i.i.d. Based on their results, using $P=O\big{(}(n\log n)/d\big{)}$ hyperplane arrangements, the stochastic objective in (3.7) attains a global minimum of zero with high probability. Consequently, standard DP-ERM bounds yield the result of Thm. 4.1. In future work, it will be interesting to see whether techniques from private linear regression (Liu et al., , 2023; Avella-Medina et al., , 2023; Varshney et al., , 2022; Cai et al., , 2021) could be used to eliminate the assumption of bounded gradients. Notice that Thm. 4.1 exhibits dimension-independence unlike existing DP-SGD bounds for the linear regression under the random data assumption (see, e.g., Thm. 1.2, Brown et al., , 2024).
@@ 4.1 Utility Bound for the Convex ReLU Approximation in the Random Data Model
Consider the random data model and consider applying the private gradient descent (repeated from literature in Alg. 1) to the strongly convex loss (3.6) and assume that the gradients are bounded by a constant $L>0$. Let the ratio $c=\frac{n}{d}\geq 1$ be fixed. For any $\gamma>0$, there exists $d_{1}$ such that for all $d\geq d_{1}$, with probability at least $1-\gamma-\frac{1}{(2n)^{8}}$ (with $\widetilde{O}$ omitting the logarithmic factors)
$\mathcal{L}(\theta^{\textrm{priv}},D)\leq\widetilde{O}\left(\frac{1}{\sqrt{n}\varepsilon}\right).$